# IoDResearch: Deep Research on Private Heterogeneous Data via the Internet of Data

### v1

@@ Abstract
The rapid growth of multi-source, heterogeneous, and multimodal scientific data has increasingly exposed the limitations of traditional data management. Most existing DeepResearch (DR) efforts focus primarily on web search while overlooking local private data. Consequently, these frameworks exhibit low retrieval efficiency for private data and fail to comply with the FAIR principles, ultimately resulting in inefficiency and limited reusability. To this end, we propose IoDResearch (Internet of Data Research), a private data-centric Deep Research framework that operationalizes the Internet of Data paradigm. IoDResearch encapsulates heterogeneous resources as FAIR-compliant digital objects, and further refines them into atomic knowledge units and knowledge graphs, forming a heterogeneous graph index for multi-granularity retrieval. On top of this representation, a multi-agent system supports both reliable question answering and structured scientific report generation. Furthermore, we establish the IoD DeepResearch Benchmark to systematically evaluate both data representation and Deep Research capabilities in IoD scenarios. Experimental results on retrieval, QA, and report-writing tasks show that IoDResearch consistently surpasses representative RAG and Deep Research baselines. Overall, IoDResearch demonstrates the feasibility of private-data-centric Deep Research under the IoD paradigm, paving the way toward more trustworthy, reusable, and automated scientific discovery.
@@ 1 Introduction
In the data-driven era of artificial intelligence, the limitations of traditional data management approaches have become increasingly evident with the explosive growth of data. Efficiently managing and leveraging massive heterogeneous data has emerged as a pressing challenge[1, 2]. This issue is particularly critical in scientific research, where building robust and efficient cross-domain data infrastructures is becoming increasingly important[3]. Real-world scientific data are often multi-source, heterogeneous, and multimodal, lacking unified representation standards and interoperability mechanisms, which leads to low efficiency in data utilization. For example, materials science research typically involves vast amounts of heterogeneous data derived from experiments, simulations, and literature records [4]. Effective management, sharing, and analysis of these data are essential for accelerating scientific discovery and fostering innovation. However, due to inherent differences in data formats, structures, and semantics, as well as the ever-increasing scale of data, traditional centralized management approaches face severe challenges [5].
In this context, the Internet of Data (IoD)[6, 7] has emerged as a data-centric paradigm designed to overcome the limitations of traditional data management. IoD leverages open, data-centric software architectures and standardized interoperability protocols to interconnect heterogeneous platforms and systems, thereby forming a unified data network. Rooted in the Digital Object Architecture (DOA)[8], IoD provides persistent identifiers, enriched metadata, and standardized interoperability mechanisms to encapsulate diverse resources as digital objects (DOs).
The FAIR principles, proposed by Wilkinson et al.[9], were established to address the lack of common standards in the face of rapidly growing scientific data. They ensure that data are Findable, Accessible, Interoperable, and Reusable. Without these four dimensions, data would remain locked in isolated systems, making cross-domain retrieval and reuse extremely difficult. By design, IoD is inherently aligned with the FAIR principles. Through these mechanisms, IoD effectively eliminates data silos, enabling seamless linking and reuse of distributed heterogeneous data, and thus lays a solid foundation for cross-domain scientific data infrastructures.
With the rapid advancement of large language models, the development of Deep Research (DR) agents[10] has gained significant attention in both academia and industry. Such agents exhibit a wide range of promising abilities, including proposing novel research directions [11, 12], efficiently acquiring information through search-augmented tools [13, 14], and performing preliminary analyses or experiments before composing full research reports or papers [15, 16].
@@
  • IoD Heterogeneous Graph Representation. We develop a three-layer IoD-based representation that transforms raw resources into FAIR-compliant digital objects, and further refines them into atomic knowledge, vector embeddings, and knowledge graphs. This enables multi-granularity indexing and retrieval, from entire objects to fine-grained facts and semantic relations.
@@ 2.2.1 IoD search tools
All retrieval functions are encapsulated as tools via MCP and exposed to agents through standardized interfaces. We distinguish retrieval granularity from retrieval strategy. In terms of granularity, we support three tiers: digital-object retrieval, which returns relevant digital objects (DOs); content-chunk retrieval, which targets passages within a DO or the content of Level-2 Digital Objects (L2-DOs); and fine-grained retrieval at the knowledge-refinement layer, which returns atomic facts or subgraphs of the knowledge graph. Independently of granularity, retrieval modules can operate under multiple strategies—including keyword search, vector-based similarity, knowledge-graph-based reasoning, and multi-source/hybrid recall. To ensure quality, each retrieved item is accompanied by metadata (e.g., type, source, timestamp), enabling agents to automatically filter outdated, unreliable, or untrusted information. This design improves retrieval robustness and enhances both interpretability and trustworthiness in downstream reasoning.
@@ 4 Conclusion
In this work, we introduced IoDResearch, a private data-centric Deep Research framework that operationalizes the IoD paradigm. By encapsulating heterogeneous assets into FAIR-compliant digital objects and refining them into atomic knowledge and knowledge graphs, IoDResearch enables high-recall and complex reasoning over private data sources. Together with a multi-agent system, it supports not only reliable question answering but also structured scientific report generation. Furthermore, we released the IoD DeepResearch Benchmark, which provides the first systematic evaluation of data representation and agentic Deep Research capabilities under the IoD paradigm. Experimental results consistently demonstrate IoDResearch’s superiority over existing RAG and DeepResearch baselines in IoD application scenarios.

### v2

@@ Abstract
The rapid growth of multi-source, heterogeneous, and multimodal scientific data has increasingly exposed the limitations of traditional data management. Most existing Deep Research (DR) efforts focus primarily on web search while overlooking local private data. Consequently, these frameworks exhibit low retrieval efficiency for private data and fail to comply with the FAIR principles, ultimately resulting in inefficiency and limited reusability. To this end, we propose IoDResearch (Internet of Data Research), a private data-centric Deep Research framework that operationalizes the Internet of Data paradigm. IoDResearch encapsulates heterogeneous resources as FAIR-compliant digital objects, and further refines them into atomic knowledge units and knowledge graphs, forming a heterogeneous graph index for multi-granularity retrieval. On top of this representation, a multi-agent system supports both reliable question answering and structured scientific report generation. Furthermore, we establish the IoD Deep Research Benchmark to systematically evaluate both data representation and Deep Research capabilities in IoD scenarios. Experimental results on retrieval, question answering, and report-writing tasks show that IoDResearch consistently surpasses representative retrieval-augmented generation (RAG) and Deep Research baselines. Overall, IoDResearch demonstrates the feasibility of private-data-centric Deep Research under the IoD paradigm, paving the way toward more trustworthy, reusable, and automated scientific discovery. URL: https://github.com/FredericVAN/PKU_IoDResearch
@@ 1 Introduction
In the data-driven era of artificial intelligence, the limitations of traditional data management approaches have become increasingly evident with the explosive growth of data, making the efficiently managing and leveraging massive heterogeneous data has emerged as a pressing challenge [16, 12], especially in scientific research scenarios that require robust cross-domain data infrastructures [7].Because Real-world scientific data are often multi-source, heterogeneous, and multimodal, lacking unified representation standards and interoperability mechanisms, which leads to low efficiency in data utilization. For example, materials science research typically involves vast amounts of heterogeneous data derived from experiments, simulations, and literature records [3]. Effective management, sharing, and analysis of these data are essential for accelerating scientific discovery and fostering innovation. However, due to inherent differences in data formats, structures, and semantics, as well as the ever-increasing scale of data, traditional centralized management approaches face severe challenges [21].
To address these limitations, the Internet of Data (IoD)[11, 20] has emerged as a data-centric paradigm designed to overcome the limitations of traditional data management. IoD leverages open, data-centric software architectures and standardized interoperability protocols to interconnect heterogeneous platforms and systems, thereby forming a unified data network. Rooted in the Digital Object Architecture (DOA)[8], IoD provides persistent identifiers, enriched metadata, and standardized interoperability mechanisms to encapsulate diverse resources as digital objects (DOs).
In parallel, the FAIR principles, proposed by Wilkinson et al.[17], were established to address the lack of common standards in the face of rapidly growing scientific data. They ensure that data are Findable, Accessible, Interoperable, and Reusable. Without these four dimensions, data would remain locked in isolated systems, making cross-domain retrieval and reuse extremely difficult. By design, IoD is inherently aligned with the FAIR principles. Through these mechanisms, IoD effectively eliminates data silos, enabling seamless linking and reuse of distributed heterogeneous data, and thus lays a solid foundation for cross-domain scientific data infrastructures.
More recently, the rapid advancement of large language models has accelerated the development of Deep Research (DR) agents[5] has gained significant attention in both academia and industry. Such agents exhibit a wide range of promising abilities, including proposing novel research directions [14, 4], efficiently acquiring information through search-augmented tools [10, 6], and performing preliminary analyses or experiments before composing full research reports or papers [18, 22].
@@
  • IoD Heterogeneous Graph Representation. We develop a three-layer IoD-based representation that transforms raw resources into FAIR-compliant digital objects(DO), and further refines them into atomic knowledge, vector embeddings, and knowledge graphs. This enables multi-granularity indexing and retrieval, from entire objects to fine-grained facts and semantic relations.
@@ 2.2.1 IoD search tools
All retrieval functions are encapsulated as tools via Model Context Protocol (MCP) and exposed to agents through standardized interfaces. We distinguish retrieval granularity from retrieval strategy. In terms of granularity, we support three tiers: digital-object retrieval, which returns relevant digital objects (DOs); content-chunk retrieval, which targets passages within a DO or the content of Level-2 Digital Objects (L2-DOs); and fine-grained retrieval at the knowledge-refinement layer, which returns atomic facts or subgraphs of the knowledge graph. Independently of granularity, retrieval modules can operate under multiple strategies—including keyword search, vector-based similarity, knowledge-graph-based reasoning, and multi-source/hybrid recall. To ensure quality, each retrieved item is accompanied by metadata (e.g., type, source, timestamp), enabling agents to automatically filter outdated, unreliable, or untrusted information. This design improves retrieval robustness and enhances both interpretability and trustworthiness in downstream reasoning.
@@ 3.2 Results and analysis
4 Limitations and Future Work
As the current pipeline is primarily text-based, information loss may occur during preprocessing and context construction for multimodal data. In future work, we will explore an end-to-end multimodal pipeline using multimodal embeddings and LLMs, and investigate post-training strategies to better assess the upper-bound performance of the proposed framework.
5 Acknowledgment
This work was supported by the National Natural Science Foundation of China under Grant No. 62595734.
@@ 6 Conclusion
In this work, we introduced IoDResearch, a private data-centric Deep Research framework that operationalizes the IoD paradigm. By encapsulating heterogeneous assets into FAIR-compliant digital objects and refining them into atomic knowledge and knowledge graphs, IoDResearch enables high-recall and complex reasoning over private data sources. Together with a multi-agent system, it supports not only reliable question answering but also structured scientific report generation. Furthermore, we released the IoD DeepResearch Benchmark, which provides the first systematic evaluation of data representation and agentic Deep Research capabilities under the IoD paradigm. Experimental results show that IoDResearch outperforms the selected RAG and Deep Research baselines on the proposed IoD DeepResearch Benchmark across the evaluated scenarios and task settings.